OpenStudy (anonymous):

expand using the properties of logarithms: log5 y(x+2)/x^4

OpenStudy (anonymous):

when expanding a logarithm when two terms are being multiplied together the sum of the logarithm of the two taken seperately is equal for example log base 5 of x(y) is equal to log base 5 of x + log base 5 of y

OpenStudy (anonymous):

when a term is being divided then you use the difference of the log with the same base for example log base 5 of x/y is equal to log base 5 of x - log base 5 of y

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

\[\log5 y(x+2)/x^4 = \log5 y(x+2) - \log5x^4 = \log5y+\log5(x+2) - 4\log5x\]

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

bubbamurphy your answer gets cut off

OpenStudy (anonymous):

Sorry the above reply got cut out. \[\log5y(x+2) - \log5x^4 = \log5y + \log5(x+2) - 4\log5x\]

OpenStudy (anonymous):

The righthand side of that last post should be your final answer (assuming by log5 you meant base 5)

OpenStudy (anonymous):

and yes then you move the power to the front

OpenStudy (anonymous):

as bubbamurphy did

OpenStudy (anonymous):

when two or multiple things are being multiplied together in the same log you can separate them with addition when they are being divided then you may use subtraction

OpenStudy (anonymous):

i did thankyou

OpenStudy (anonymous):

-4log5x is my answer?

OpenStudy (anonymous):

the whole left side of the equation that bubbamurphy posted is your completed answer

OpenStudy (anonymous):

*right side

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

no thankyou allot

OpenStudy (anonymous):

no problem any other questions while i am here?

OpenStudy (anonymous):

expand the expression: log3(x^-2y^3)

OpenStudy (anonymous):

\[\log3(x^(-2y^3) = -2y^3\log3x\]

OpenStudy (anonymous):

what is above the first x?

OpenStudy (anonymous):

the way that you wrote the problem x was to the power of -2y^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if x is to the power of that entire quantity you can move that quantity to the front of the logarithm

OpenStudy (anonymous):

quantity meaning -2y^3

OpenStudy (anonymous):

that is a parenthesis above the first x

OpenStudy (anonymous):

there should be a end parenthesis after the -2y^3

OpenStudy (anonymous):

thankyou for your help

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

do you have time for more?